The product of two consecutive positive integers is 56. What is the smaller integer?
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

Using some guessing and checking, we see that the integers are 7 and 8.

Here is a more algebraic approach:

smaller integer = x

x(x+1) = 56
x^2 + x = 56
x^2 + x - 56 = 0
(x + 8)(x - 7) = 0

x is positive, so x = 7.

Answer by ikleyn(52790)

.

This problem/question is to check if you know the multiplication table . . .